# Best Ergodic Averages via Optimal Graph Filters in Reversible Markov Chains

### v1

@@ Abstract
In this paper, we address the problem of finding the best ergodic or Birkhoff averages in the ergodic theorem to ensure rapid convergence to a desired value, using graph filters. Our approach begins by representing a function on the state space as a graph signal, where the (directed) graph is formed by the transition probabilities of a reversible Markov chain. We introduce a concept of graph variation, enabling the definition of the graph Fourier transform for graph signals on this directed graph. Viewing the iteration in the ergodic theorem as a graph filter, we recognize its non-optimality and propose three optimization problems aimed at determining optimal graph filters. These optimization problems yield the Bernstein, Chebyshev, and Legendre filters. Numerical testing reveals that while the Bernstein filter performs slightly better than the traditional ergodic average, the Chebyshev and Legendre filters significantly outperform the ergodic average, demonstrating rapid convergence to the desired value.
Index Terms:
@@
Ergodic theorem, graph filters, Bernstein-Chebyshev-Legendre polynomials.
I Introduction
@@
The ergodic theorem is a cornerstone in the study of dynamical systems and stochastic processes, providing a foundation for understanding long-term average behavior. Traditional ergodic or Birkhoff averages are essential tools in this significant result, but their convergence rates can be slow. This paper seeks to enhance these averages using graph signal processing techniques, particularly through the application of graph filters.
@@
Our approach begins by conceptualizing a function on the state space as a graph signal, where the (directed) graph, on which a graph signal can be defined, is constructed based on the transition probabilities of a reversible Markov chain. By introducing a notion of graph variation, we can define the graph Fourier transform for graph signals on this directed graph. This new perspective allows us to interpret the iteration in the ergodic theorem as a graph filter. However, this standard ergodic graph filter is not optimal in terms of convergence speed. To address this issue, we formulate three distinct optimization problems, each with a different objective, to derive the optimal graph filters. The solutions to these problems result in the Bernstein, Chebyshev, and Legendre polynomial filters.
@@
We then validate our theoretical findings through numerical experiments. The results indicate that the Bernstein filter offers a slight improvement over the ergodic average. More notably, the Chebyshev and Legendre filters exhibit significantly better performance, achieving rapid convergence to the desired value. These findings highlight the effectiveness of our proposed graph signal processing approach in optimizing ergodic averages.
Our paper intersects with two main lines of research: faster algorithm design via approximation theory and optimal Birkhoff averages in ergodic theorem for deterministic dynamical systems. The first domain aims to accelerate algorithms by approximating some well-known functions, which appear in the computation of outputs of these algorithms, using polynomials [1]. The second focuses on selecting optimal ergodic averages to enhance the convergence rates in ergodic theorem for deterministic dynamical systems [2].
In algorithm design, approximation theory plays a crucial role in developing efficient computational methods. Early applications include the conjugate gradient method [3, 4] and the Lanczos method [5, 6], both of which are used to solve systems of linear equations. These techniques significantly reduce the number of computations compared to traditional methods like Gaussian elimination. Furthermore, approximation theory also extends its usage to complexity theory. For instance, Chebyshev polynomials have been used to enhance quantum search algorithms [7]. Particularly relevant to our study is the relatively recent monograph [1], which explores how approximating well-known functions with polynomials or rational functions can accelerate algorithms. This monograph studies techniques for approximating functions such as $x^{s}$, $x^{-1}$, and $e^{-x}$ and investigates how these results play an important role in the design of fast algorithms for some well-known problems. In particular, these techniques facilitate more efficient computation of fundamental objects like $A^{s}v$, $A^{-1}v$, $\exp(-A)v$, eigenvalues, and eigenvectors, which are integral to many spectral algorithms. In fact, numerous algorithms can be reduced to the calculation of these objects, which have been shown to significantly expedite essential computations such as random walk simulations, graph partitioning, and solving systems of linear equations.
In this monograph, the primary objective is to approximate the aforementioned functions with low-degree polynomials given a specified maximum error, without an optimization component. In contrast, our approximations are performed optimally by formulating three optimization problems with distinct objectives to identify the best polynomial, resulting in three well-known polynomials in approximation theory. This approach allows practitioners to select the most suitable polynomial based on the specific application by considering the appropriate objective. Furthermore, while the monograph deals with fixed functions, our approach specifies the function to approximate only over a sub-interval of its domain and on a single isolated point. This distinction makes our problem significantly different from those addressed in the monograph. Additionally, our paper introduces a novel perspective to these approximation problems by interpreting them as graph filters. This enables us to use established signal processing methods, a viewpoint that is absent in the monograph.
@@
In the context of ergodic theory for deterministic dynamical systems, the Birkhoff ergodic theorem states that, for ergodic systems, the time averages of a function along a trajectory of length $T$ will converge to the space average as $T$ approaches infinity. However, this convergence is often slow. To address this, researchers have explored using non-uniform weights for averages, instead of uniform averages that equally weight each point along the trajectory. This method is particularly relevant to our problem formulation. In [2], the authors address this issue by assigning significantly lower weights to the initial and final points of the trajectory compared to those near the midpoint using a bump function. They demonstrate that, in quasiperiodic dynamical systems, these weighted averages achieve much faster convergence, provided that the function $f$ has sufficient differentiability. The studies in [8, 9, 10] also discuss the problem of accelerating the convergence of weighted ergodic sums, yet they do not specify the exact rates of convergence as presented in [2]. In [11], for a certain choice of weights, a convergence rate is obtained which is worse than the one in [2]. In [12], authors established the polynomial and exponential convergence rate of weighted Birkhoff averages of irrational rotations on tori, which improves the result in [2]. In these works, there is no optimization component in the problem formulation as opposed to our case. Namely, it is established that if the function generating the weights satisfies certain properties, the weighted Birkhoff averages converge to the space average at some speed. However, they do not claim that this is the optimal convergence rate achievable, nor do they assert that the weights are optimal.
@@
In this direction, the most pertinent work related to our study is [13]. In this work, the authors expand the function $f$, whose integral is to be computed, using basis functions and view the weighted Birkhoff averages as a filter on the coefficients of this expansion. They observe that a filter tuned to specific frequencies can be significantly more effective than an arbitrary bump function, as considered in [2]. This approach aligns with the method we take in this paper. However, there are notable distinctions. First, [13] considers deterministic dynamical systems on abstract topological spaces, whereas we focus on stochastic dynamics given by a Markov chain on a finite set. Additionally, in [13], there are infinitely many frequency components, and so, the function approximated via filters differs from ours. Furthermore, the basis functions corresponding to the frequencies are not related to the eigenfunctions of the dynamical map, meaning there is no operator-theoretic perspective or spectral analysis of the dynamical map. Finally, there is no underlying graph structure in their work, unlike in our case. It is worth mentioning that our work was developed independently of [13]. Consequently, our motivation for employing a filtering approach to achieve fast convergence in the ergodic theorem for Markov chains stems entirely from graph signal processing theory. Nevertheless, it is fascinating to see that such a natural approach is also being embraced by the dynamical systems community.
@@
We note that even if one views optimal Birkhoff averages in deterministic dynamical systems via an operator theory perspective, a complication still arises as the operators are generally infinite-dimensional, and so their spectra are not well-behaved (isolated) as in the finite-dimensional case. Hence, graph signal processing techniques cannot be applied. Indeed, an important research problem in this direction is as follows: consider a Markov chain with an abstract state space. Then the transition probability can be viewed as an infinite-dimensional operator, where the spectrum may contain infinitely many elements. In this case, the eigenvalue $1$ is probably not isolated, and so we cannot directly apply graph filters as we cannot easily attenuate eigenvalues that are closer to $1$. In this case, we can first approximate the operator with a finite-dimensional one and then apply our graph filters. This is indeed a research direction we want to pursue in the future.
This paper is structured as follows: Section II reviews reversible Markov chains. Section III details our perspective about the ergodic theorem, including the construction of the graph signal framework and the definition of graph Fourier transform. Sections IV, V, and VI present the formulations and solutions of the optimization problems that lead to Bernstein, Chebyshev, and Legendre polynomial filters, respectively. Section VII discusses our numerical experiments and results. Finally, Section VIII offers concluding remarks.
II Reversible Markov Chains
@@
Let ${\mathsf{X}}$ be a finite set and let $P:{\mathsf{X}}\rightarrow{\mathcal{P}}({\mathsf{X}})$ be a transition probability of a Markov chain on ${\mathsf{X}}$. By an abuse of notation, we also denote the transition matrix via $P$, and so, $P\in\mathbb{R}^{{\mathsf{X}}\times{\mathsf{X}}}$. Therefore, $P(x,y)=\mathop{\rm Pr}\{X_{t+1}=y|X_{t}=x\}$ for all $x,y\in{\mathsf{X}}$. We assume that $P$ is irreducible. In this case, it is known that there exists a unique stationary distribution $\pi$ of $P$ (see [14, Corollary 1.17]) and $\pi(x)>0$ for all $x\in{\mathsf{X}}$ (see [14, Proposition 1.19]).
Definition 1.
The transition probability $P$ is said to be reversible if there exists a probability distribution $\pi$ on ${\mathsf{X}}$ such that
$\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\,\,\forall x,y\in{\mathsf{X}}.$
The above equation is called detailed balance equation.
By [14, Proposition 1.20], any probability distribution that satisfies the detailed balance equation is a stationary distribution of $P$:
$\big{[}\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\,\,\forall x,y\in{\mathsf{X}}\big{]}\implies\big{[}\pi P=\pi\big{]}.$
Moreover, the corresponding Markov chain $\{X_{t}\}$ with initial distribution $\pi$ satisfies the following: for any $n$, we have
$\mathop{\rm Pr}\{X_{0}=x_{0},X_{1}=x_{1},\ldots,X_{n}=x_{n}\}$
$=\pi(x_{0})P(x_{0},x_{1})\ldots P(x_{n-1},x_{n})$
$=\pi(x_{n})P(x_{n},x_{n-1})\ldots P(x_{1},x_{0})$
$=\mathop{\rm Pr}\{X_{0}=x_{n},X_{1}=x_{n-1},\ldots,X_{n}=x_{0}\}.$
@@
Hence, the time reversed version of the Markov chain $\{X_{t}\}$ is indistinguishable from the original Markov chain.
Example 1.
Random Walk on Graphs
Since the graph is connected, the corresponding Markov chain is irreducible. It is also straightforward to see that the Markov chain is reversible with respect to
$\pi(x)=\frac{d(x)}{2|E|},\,\,x\in{\mathsf{X}},$
where $|E|$ is the number of edges.
Example 2.
Glauber Chains
In this example, $V$ and $S$ are finite sets. We can think of $V$ as the set of vertices of a graph and $S$ is the set of possible values that can be taken by these vertices. Our state space ${\mathsf{X}}$ is a subset of $V^{S}$. Let $\pi$ be a probability distribution on ${\mathsf{X}}$. In statistical physics, $\pi$ is called a Gibbs distribution.
Here, ${\mathsf{X}}(x,v)\coloneqq\{y\in{\mathsf{X}}:y(w)=x(w)\,\,\forall\,w\neq v\}$. Verbally, a new state is chosen according to the probability distribution $\pi$ conditioned on the set of states equal to the current state $x$ at all vertices different from $v$.
Define ${\mathsf{X}}(x)\coloneqq\bigcup_{v\in V}{\mathsf{X}}(x,v)$. Then, the transition probability can be defined as follows:
A specific Glauber dynamics can be constructed if $\pi$ is a Gibbs distribution of the following form:
$\pi(x)=\frac{e^{-\beta H(x)}}{Z(\beta)},$
where $H(x)$ is the energy of the state $x$ and $Z(\beta)\coloneqq\sum_{x\in{\mathsf{X}}}e^{-\beta H(x)}$ is the partition function. For instance, if $S=\{-1,+1\}$ and $V$ is the vertex set of some finite graph, then
$H(x)=-\sum_{u,w\in V:u\sim w}J_{v,w}\,x(v)x(w),$
where $J_{v,w}>0$. Then, the transition probability is given by
[…]
where $S(x,w)\coloneqq\sum_{u:u\sim w}J_{u,w}\,x(u)$.
II-A Eigenvalues of Reversible Markov Chains
Let $P$ be a reversible Markov chain with respect to $\pi$. Then, we define the following inner product on $\mathbb{R}^{{\mathsf{X}}}$:
$\langle f,g\rangle_{\pi}\coloneqq\sum_{x\in{\mathsf{X}}}f(x)g(x)\pi(x).$
The following result follows from [14, Lemmas 12.1 and 12.2].
Lemma 1.
@@ II-A Eigenvalues of Reversible Markov Chains
    There exists orthonormal basis of real-valued eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$ corresponding to real eigenvalues $\{\lambda_{x}\}_{x\in{\mathsf{X}}}$ for inner product space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$.
The spectral gap of a reversible Markov chain $P$ is defined as $\gamma\coloneqq 1-\lambda_{2}$, where $\lambda_{2}$ is the second largest eigenvalue of $P$. The spectral gap, or a lower bound to it, is important in the optimal graph filter design. We close this subsection by stating the Birkhoff ergodic theorem for irreducible Markov chains.
Theorem 1.
Birkhoff Ergodic Theorem
@@
Hence, ergodic theorem implies that $\frac{1}{t}(\mathbb{Id}+P+\ldots+P^{t-1})f$ converges to the constant vector $\left(\sum_{z\in{\mathsf{X}}}f(z)\pi(z)\right){\bf 1}$ pointwise. Let $\pi(f)=\sum_{z\in{\mathsf{X}}}f(z)\pi(z)$. The iteration in ergodic theorem
@@
can be considered as a low-pass graph filter for the graph signal $f$, as we will elaborate on in the next section. Therefore, in the remainder of this paper, we aim to develop algorithms to identify the optimal low-pass graph filter that achieves convergence to $\pi(f){\bf 1}$ faster than the aforementioned iteration.
III Graph Signal Processing
In this section, ${\mathcal{G}}=({\mathsf{X}},E)$ denotes a simple (directed) graph with a finite vertex set ${\mathsf{X}}$ and an edge set $E$. Therefore, if $(x,y)\in E$, there is a directed edge from $x$ to $y$, denoted as $x\rightarrow y$. Each directed edge in this graph has an associated weight, denoted by $P(x,y)\in\mathbb{R}_{+}$. Since we are interested in Markov chains, we assume that for any $x$, $\sum_{y\in{\mathsf{X}}}P(x,y)=1$; in other words, $P$ can be viewed as the transition probability of a Markov chain. Additionally, there is a probability distribution $\pi$ on ${\mathsf{X}}$ such that the weight matrix $P$ satisfies the detailed balance equation:
$\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\forall x,y\in{\mathsf{X}};$
that is, $P$ is reversible. Hence, although graph signal processing techniques can be applied to any directed and weighted graph, in the remainder of this paper, we assume that the weight matrix $P$ corresponds to the transition probability of a reversible and irreducible Markov chain with respect to the unique stationary distribution $\pi$. Moreover, we endow $\mathbb{R}^{{\mathsf{X}}}$ with the inner product $\langle\cdot,\cdot\rangle_{\pi}$ introduced in the previous section. We direct the reader to the book [15] for the terminology and notation related to graph signal processing used in this paper.
In graph ${\mathcal{G}}$, for any vertex $x$, its out-degree is $d_{out}(x)=\sum_{x\rightarrow y}P(x,y)=1$. Consequently, the out-degree matrix for this graph is the identity matrix $\mathbb{Id}$. Therefore, the combinatorial graph Laplacian is defined as $L\coloneqq\mathbb{Id}-P$. Given that $\sigma(P)\subset[-1,1]$, it follows $\sigma(L)\subset[0,2]$.
Remark 1.
Given that $P$ satisfies the detailed balance equation with respect to $\pi$, we can define the following symmetric matrix:
$A(x,y)\coloneqq\pi(x)^{1/2}P(x,y)\pi(y)^{-1/2}\,\,\forall x,y\in{\mathsf{X}}.$
In matrix notation, if we let $D\coloneqq\mathop{\rm diag}(\pi)$ (i.e., the diagonal matrix with diagonal entries formed by the vector $\pi$), then, $A=D^{1/2}PD^{-1/2}$. Hence, $A$ and $P$ are similar matrices. Similar matrices represent the same linear operator in different bases, and they share all properties of the underlying operator, including the identical set of (real) eigenvalues. Consequently, we can use $A$ as a weight matrix instead of $P$. In this case, since $A$ is symmetric, we can treat our graph as a weighted undirected graph. However, a complication arises because the stationary distribution $\pi$ is either unknown or difficult to compute precisely, making it challenging to construct $A$ for use in graph signal processing.

### v2

@@ Abstract
In this paper, we address the problem of finding the best ergodic or Birkhoff averages in the mean ergodic theorem to ensure rapid convergence to a desired value, using graph filters. Our approach begins by representing a function on the state space as a graph signal, where the (directed) graph is formed by the transition probabilities of a reversible Markov chain. We introduce a concept of graph variation, enabling the definition of the graph Fourier transform for graph signals on this directed graph. Viewing the iteration in the mean ergodic theorem as a graph filter, we recognize its non-optimality and propose three optimization problems aimed at determining optimal graph filters. These optimization problems yield the Bernstein, Chebyshev, and Legendre filters. Numerical testing reveals that while the Bernstein filter performs slightly better than the traditional ergodic average, the Chebyshev and Legendre filters significantly outperform the ergodic average, demonstrating rapid convergence to the desired value.
keywords:
@@
Mean ergodic theorem, graph filters, Bernstein-Chebyshev-Legendre polynomials.
{MSCcodes}
94A12,41A10, 05C50, 65D15
1 Introduction
@@
The (mean) ergodic theorem is a cornerstone in the study of dynamical systems and stochastic processes, providing a foundation for understanding long-term average behavior. Traditional (mean) ergodic or Birkhoff averages are essential tools in this significant result, but their convergence rates can be slow. This paper seeks to enhance these averages using graph signal processing techniques, particularly through the application of graph filters.
@@
Our approach begins by conceptualizing a function on the state space as a graph signal, where the (directed) graph, on which a graph signal can be defined, is constructed based on the transition probabilities of a reversible Markov chain. By introducing a notion of graph variation, we can define the graph Fourier transform for graph signals on this directed graph. This new perspective allows us to interpret the iteration in the mean ergodic theorem as a graph filter. However, this graph filter is not optimal in terms of convergence speed. To address this issue, we formulate three distinct optimization problems, each with a different objective, to derive the optimal graph filters. The solutions to these problems result in the Bernstein, Chebyshev, and Legendre polynomial filters.
@@
We then validate our theoretical findings through numerical experiments. The results indicate that the Bernstein filter offers a slight improvement over the standard ergodic average. More notably, the Chebyshev and Legendre filters exhibit significantly better performance, achieving rapid convergence to the desired value. These findings highlight the effectiveness of our proposed graph signal processing approach in optimizing mean ergodic averages.
Our paper lies at the intersection of two major lines of research: acceleration techniques in numerical linear algebra based on approximation theory, and the study of optimal Birkhoff averages in the ergodic theory of deterministic dynamical systems. The first line seeks to speed up iterative algorithms for numerical linear algebra problems by designing polynomial filters that selectively attenuate undesirable spectral components [17]. The second investigates how to construct optimal ergodic averages that improve convergence rates in the ergodic theorem for deterministic systems [5].
In numerical linear algebra tasks such as solving linear systems or computing extreme eigenvalues are frequently accelerated using polynomial filtering techniques, either adaptive or non-adaptive. Classical examples include the Lanczos method [9, 18], which can be viewed as an adaptive polynomial filtering procedure for computing extreme eigenvalues of symmetric matrices, and its more general counterpart, the Arnoldi iteration [1]. In these methods, convergence is enhanced by adaptively reshaping the spectrum of the underlying matrix. Chebyshev acceleration, also known as the Chebyshev semi-iterative method [21, 12], provides a complementary, non-adaptive approach, constructing polynomial filters that suppress unwanted spectral components while amplifying the desired ones [17]. Indeed, the optimization problem that gives rise to the Chebyshev filter in our setting coincides with the one underlying Chebyshev acceleration (or the Chebyshev semi-iterative method). This correspondence is entirely natural, as the construction of near-optimal polynomial filters through approximation-theoretic techniques is a classical problem that has been extensively developed in numerical analysis and signal processing. Accordingly, we do not claim novelty in this approximation-theoretic aspect, as in this part we simply apply well-established results from approximation theory within our problem framework. Our contribution instead lies in introducing a new perspective, namely the interpretation of the ergodic theorem through the framework of graph signal processing.
Specifically, we formulate the acceleration of ergodic averages as a graph signal processing problem. Functions acted upon by a Markov kernel are interpreted as graph signals on the directed graph induced by the transition probabilities. Within this framework, we introduce a notion of graph variation, relate it to the Laplacian spectrum of the graph, and interpret the Laplacian eigenvalues as frequencies. Under this viewpoint, ergodic convergence becomes a low-pass filtering problem: the zero Laplacian eigenvalue represents the lowest frequency, and the associated Fourier coefficient corresponds to the expectation with respect to the stationary distribution. The goal is therefore to eliminate higher-frequency components and retain only the zero-frequency mode.
This objective differs fundamentally from that of eigenvalue algorithms such as Lanczos and Arnoldi. Those methods are designed to compute extreme eigenvalues and their corresponding eigenvectors. In our setting, however, these quantities are already known: the smallest Laplacian eigenvalue is zero, and its eigenvector is the constant vector $\mathbf{1}$. Our aim is therefore not to compute this eigenpair, but rather to suppress all other spectral components of a given signal in order to extract its stationary mean.
There is also a structural distinction. Arnoldi-type methods are adaptive and require repeated orthogonalization, leading to substantial computational and memory costs. In contrast, our approach employs non-adaptive polynomial filters, similar in spirit to Chebyshev semi-iterative method. The filter is fixed in advance, depends only on spectral bounds, and can be implemented through simple iterative recursions that preserve the lightweight structure of the classical Birkhoff ergodic iteration. Indeed, as our goal is to generalize that iteration while maintaining its simplicity, adaptive procedures are not suitable for our purposes.
In summary, although polynomial spectral filtering is a classical technique and closely related to acceleration methods in numerical linear algebra, our contribution lies in the novel graph signal processing interpretation of ergodic convergence. By formulating ergodic acceleration as a low-pass filter design problem tailored to reversible Markov chains, we provide a fresh perspective that opens up new research directions. This viewpoint suggests that additional filtering techniques from the signal processing literature, such as IIR filters, could potentially be adapted to this setting to further enhance convergence. More broadly, framing the problem within signal processing allows us to draw upon a rich set of tools and methodologies from that field to address related challenges. We believe this perspective will inspire further work at the intersection of dynamical systems, Markov chains, and graph signal processing.
In the context of ergodic theory for deterministic dynamical systems, the Birkhoff ergodic theorem states that, for ergodic systems, the time averages of a function along a trajectory of length $T$ converge to the corresponding space average as $T\to\infty$. In practice, however, this convergence is often slow. To address this limitation, researchers have explored the use of non-uniform weights in place of the standard uniform averages that weight each point along the trajectory equally; an approach that is particularly relevant to our problem formulation.
@@
In [5], the authors assign significantly lower weights to the initial and final points of the trajectory relative to those near the midpoint using a bump function. They demonstrate that, for quasi-periodic dynamical systems, these weighted averages achieve substantially faster convergence, provided the observable $f$ is sufficiently differentiable. Related studies [11, 3, 2] also investigate the acceleration of weighted ergodic sums, though they do not provide explicit convergence rates as in [5]. In [7], a specific choice of weights yields a convergence rate that is inferior to that obtained in [5]. More recently, [19] establishes polynomial and exponential convergence rates for weighted Birkhoff averages of irrational rotations on tori, thereby improving upon the results in [5]. Importantly, none of these works incorporate an optimization component into the problem formulation. They establish that if the weight-generating function satisfies certain properties, then the weighted averages converge to the space average at a certain speed; however, they do not claim optimality of the convergence rate nor that the chosen weights are optimal in any sense.
@@
The most pertinent work to our study in this direction is [16]. There, the authors expand the observable $f$ in a basis and interpret weighted Birkhoff averages as a filter applied to the expansion coefficients. They observe that a filter tuned to specific frequencies can significantly outperform an arbitrary bump function of the type considered in [5]. This perspective aligns closely with the approach we adopt in the present paper. Nevertheless, several important distinctions remain. First, [16] considers deterministic dynamical systems on abstract topological spaces, whereas our focus is on stochastic dynamics governed by a Markov chain on a finite state space. Second, in [16], the expansion involves infinitely many frequency components, in contrast to the finite spectral decomposition that arises in our setting; consequently, the function approximated via filters differs, as infinitely many frequency components may contribute. Third, the basis functions corresponding to the frequencies in [16] are not related to the eigenfunctions of the dynamical map, and thus no operator-theoretic or spectral analysis of the dynamics is employed. Finally, their work does not involve an underlying graph structure, whereas our formulation explicitly leverages the graph induced by the Markov transition kernel.
@@
We note that even if one views optimal Birkhoff averages in deterministic dynamical systems via an operator theory perspective, a complication still arises as the operators are generally infinite-dimensional, and so their spectra are not well-behaved (isolated) as in the finite-dimensional case. Hence, graph signal processing techniques cannot be applied. Indeed, an important research problem in this direction is as follows: consider a Markov chain with an abstract state space. Then the transition probability can be viewed as an infinite-dimensional operator, where the spectrum may contain infinitely many elements. In this case, the eigenvalue $1$ is probably not isolated, and so we cannot directly apply graph filters as we cannot easily attenuate eigenvalues that are closer to $1$. In this case, we can first approximate the operator with a finite-dimensional one and then apply our graph filters.
1.1 Our Contributions
Standard ergodic averages are essential in mean ergodic theorem for Markov chains but they often suffer from slow convergence rates. This paper proposes a method to improve these averages by employing graph signal processing techniques, particularly through the use of graph filters.
  • (a)
    In Section 3, we initiate our method by interpreting a function on the state space as a graph signal. This graph signal is defined on a directed graph constructed from the transition probabilities of a reversible Markov chain. Introducing the concept of graph variation enables us to establish the graph Fourier transform for these signals on the directed graph. This framework allows us to view the iteration in the mean ergodic theorem as a graph filter. However, this graph filter does not achieve optimal convergence speed.
  • (b)
    To tackle the problem of suboptimal convergence speed, we introduce three distinct optimization problems in Sections 4, 5, and 6. Each optimization problem is designed with a unique objective in mind, aimed at deriving the most efficient graph filters for our purposes. The outcomes of these optimizations are the Bernstein, Chebyshev, and Legendre polynomial filters. By solving these optimization problems, we develop a set of graph filters that significantly enhance (at least for Chebyshev and Legendre polynomial filters) the convergence rates of mean ergodic averages.
  • (c)
    In Section 7, we support our theoretical developments with numerical experiments. The results demonstrate that while the Bernstein filter provides a modest enhancement over the traditional ergodic average, the Chebyshev and Legendre filters show substantial performance improvements.
2 Reversible Markov Chains
@@
Let $\mathsf{X}$} be a finite set and let $P:{\mathsf{X}}\rightarrow{\mathcal{P}}({\mathsf{X}})$ be a transition probability of a Markov chain on $\mathsf{X}$}. By an abuse of notation, we also denote the transition matrix via $P$, and so, $P\in\mathbb{R}^{{\mathsf{X}}\times{\mathsf{X}}}$. Therefore, $P(x,y)=\mathop{\rm Pr}\{X_{t+1}=y|X_{t}=x\}$ for all $x,y\in{\mathsf{X}}$. We assume that $P$ is irreducible. In this case, it is known that there exists a unique stationary distribution $\pi$ of $P$ (see [10, Corollary 1.17]) and $\pi(x)>0$ for all $x\in{\mathsf{X}}$ (see [10, Proposition 1.19]). The transition probability $P$ is said to be reversible if there exists a probability distribution $\pi$ on $\mathsf{X}$} such that $\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\,\,\forall x,y\in{\mathsf{X}}.$ This equation is called detailed balance equation. By [10, Proposition 1.20], any probability distribution that satisfies the detailed balance equation is a stationary distribution of $P$. Moreover, the corresponding Markov chain $\{X_{t}\}$ with initial distribution $\pi$ satisfies the following: for any $n$, we have
$\mathop{\rm Pr}\{X_{0}=x_{0},X_{1}=x_{1},\ldots,X_{n}=x_{n}\}=\pi(x_{0})P(x_{0},x_{1})\ldots P(x_{n-1},x_{n})$
$=\pi(x_{n})P(x_{n},x_{n-1})\ldots P(x_{1},x_{0})=\mathop{\rm Pr}\{X_{0}=x_{n},X_{1}=x_{n-1},\ldots,X_{n}=x_{0}\}.$
@@
Hence, the time-reversed version of the Markov chain $\{X_{t}\}$ is indistinguishable from the original chain, hence the term reversible Markov chain. We define the following inner product on $\mathbb{R}^{{\mathsf{X}}}$:
$\langle f,g\rangle_{\pi}\triangleq\sum_{x\in{\mathsf{X}}}f(x)g(x)\pi(x).$
The following result follows from [10, Lemmas 12.1 and 12.2].
Lemma 2.1.
@@
    There exists orthonormal basis of real-valued eigenfunctions $\{f_{x}\}_{x\in{\mathsf{X}}}$ corresponding to real eigenvalues $\{\gamma_{x}\}_{x\in{\mathsf{X}}}$ for inner product space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$. Moreover, for $\gamma_{x^{*}}=1$, we have $f_{x^{*}}=\mathbf{1}$.
The spectral gap of a reversible Markov chain $P$ is defined as $\gamma\triangleq 1-\lambda_{2}$, where $\lambda_{2}$ is the second largest eigenvalue of $P$. We close this section by stating the Birkhoff ergodic theorem for irreducible Markov chains.
Theorem 2.2.
@@
Hence, ergodic theorem implies that $\frac{1}{t}(\mathbb{Id}+P+\ldots+P^{t-1})f$ converges to the constant vector $\left(\sum_{z\in{\mathsf{X}}}f(z)\pi(z)\right){\bf 1}$ pointwise (or with respect to the norm induced by the inner product $\langle\cdot,\cdot\rangle_{\pi}$).
Let $\pi(f)=\sum_{z\in{\mathsf{X}}}f(z)\pi(z)$. From a graph signal processing perspective, any function on $\mathsf{X}$} is viewed as a graph signal residing in the space $(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle_{\pi})$, where the directed graph structure is induced by the transition probability $P$. The Graph Fourier Transform (GFT) is then defined as
$\hat{f}(x)=\langle f,f_{x}\rangle_{\pi}\in(\mathbb{R}^{{\mathsf{X}}},\langle\cdot,\cdot\rangle),$
where $\{f_{x}\}_{x\in{\mathsf{X}}}$ are the eigenvectors of the Laplacian, with eigenvalues $\{\lambda_{x}\}_{x\in{\mathsf{X}}}\subset[0,2]$.
[…]
Figure 1: Graph Fourier Transform
A large eigenvalue $\lambda_{x}$ indicates strong signal variation (high frequency), whereas a small $\lambda_{x}$ (low frequency) corresponds to low variation (see Lemma 3.3). In the extreme case where $\lambda_{x^{*}}=0$, we have $f_{x^{*}}=\mathbf{1}$, representing no variation at all. Since our goal is to compute $\pi(f)=\langle f,f_{x^{*}}\rangle_{\pi}$, we should attenuate all high-frequency components except for $\lambda_{x^{*}}=0$; in other words, we need to design a low-pass filter.
[…]
Figure 2: Low Pass Filter
The iteration in the ergodic theorem,
@@
can then be interpreted as a low-pass polynomial graph filter applied to the graph signal $f$. However, this particular polynomial filter is not optimal with respect to certain well-known objective functions. Motivated by this limitation, and following the classical signal processing literature [8], this paper develops algorithms to identify the optimal low-pass polynomial graph filter; one that achieves faster convergence to $\pi(f)\mathbf{1}$ than the standard ergodic average. The mathematical formulation of this approach is presented in the next section.
3 Graph Signal Processing Approach to Birkhoff Ergodic Theorem
In this section, ${\mathcal{G}}=({\mathsf{X}},E)$ denotes a directed graph with a vertex set $\mathsf{X}$} (i.e. state space) and an edge set $E$. Therefore, if $(x,y)\in E$, there is a directed edge from $x$ to $y$, denoted as $x\rightarrow y$. Each directed edge in this graph has an associated weight given by our reversible (with respect to the distribution $\pi$) transition probability $P(x,y)\in\mathbb{R}_{+}$. Hence, $P$ satisfies the detailed balance equation: $\pi(x)P(x,y)=\pi(y)P(y,x)\,\,\forall x,y\in{\mathsf{X}}.$ In view of this construction, we endow $\mathbb{R}^{{\mathsf{X}}}$ with the inner product $\langle\cdot,\cdot\rangle_{\pi}$ introduced in the previous section. We direct the reader to the book [14] for the terminology and notation related to graph signal processing used in this paper.
In graph $\mathcal{G}$}, for any vertex $x$, its out-degree is $d_{out}(x)=\sum_{x\rightarrow y}P(x,y)=1$. Consequently, the out-degree matrix for this graph is the identity matrix $\mathbb{Id}$. Therefore, the combinatorial graph Laplacian is defined as $L\triangleq\mathbb{Id}-P$. Given that $\sigma(P)\subset[-1,1]$, it follows $\sigma(L)\subset[0,2]$.